OpenStudy (anonymous):

divide 4 5/8 divided by 4 5/6

OpenStudy (anonymous):

change to improper fractions: 37/8 / 29/6 = 37/8 * 6/29 = 37/4 * 3/29 = 111/ 116

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

this question looks familiar ...